# Towards AI Search Paradigm

### v1

@@ 4.2.1 LLM Labeling
Compared to manual labeling processes, LLMs can efficiently annotate vast datasets at scale, thereby substantially reducing both the temporal and human resource investments required for training retrieval and ranking models. Furthermore, LLMs deliver highly consistent annotations, mitigating the subjective biases and inter-annotator variability that commonly afflict human-generated labels. By adjusting prompt designs or fine-tuning for specific domains or tasks, LLMs exhibit exceptional flexibility in meeting diverse annotation requirements. As previously noted, the AI search system utilizes listwise methods for document annotation, as these approaches effectively strike a balance between effectiveness and efficiency. In this work, the AI search system primarily utilizes two methods as follows.: RankGPT (Sun et al., 2023b) and TourRank (Chen et al., 2025a).
RankGPT. When LLMs come to ranking tasks, especially in the context of passage re-ranking, the limited context window of LLMs poses a significant challenge. Most LLMs have a maximum context length that restricts the number of passages they can process simultaneously. This limitation is particularly problematic in scenarios where a large number of candidate documents need to be ranked, such as in web search or large-scale IR systems. To address this challenge, the AI search system utilizes a strategy called the sliding window approach RankGPT (Sun et al., 2023b). This method allows the efficient re-ranking of large candidate sets without exceeding the context window limitations of the LLM. By breaking down the ranking task into smaller, manageable segments, RankGPT ensures that the LLM can effectively process and rank each subset of documents while maintaining the overall ranking quality.
@@
  • Initialization. Given the retrieved set of $M$ candidate documents from the retrieval model, the sliding window strategy begins with the last $w$ documents, i.e., from the $(M-w+1)$-th to the $M$-th passage. These documents are input to the LLM, which returns a ranked order for this subset. The choice of $w$ (window size) is crucial as it determines the number of documents processed in each iteration. A larger window size allows the LLM to consider more documents at once, potentially improving the ranking quality, but also increases the computational load and the risk of exceeding the context window limitations.
  • Sliding. Once the initial window is ranked, the window slides backward by $s$ documents. The new window now covers documents from $(M-w-s+1)$ to $(M-s)$. The LLM is again used to rank this subset, which may overlap with the previous window. The step size $s$ controls the overlap between consecutive windows. A smaller step size increases the overlap, allowing the LLM to compare more documents across windows, but also increases the number of iterations required to cover all documents.
@@
TourRank. LLMs have shown great potential in zero-shot document ranking. However, existing LLM-based ranking methods face several limitations. First, the maximum context length of LLMs restricts the number of documents that can be processed simultaneously, making it challenging to rank a large number of documents at once. Second, the ranking results are highly sensitive to the order in which documents are presented to the LLM, leading to inconsistent outcomes. Third, achieving a balance between computational cost and ranking performance remains a significant challenge. To address these issues, the AI search system introduces a novel approach inspired by the structure of sports tournaments, named TourRank (Chen et al., 2025a).
@@ 4.2.4 Distillation of LLM Ranking
Consider a query $q$ and its corresponding final ranking of candidate documents $\{\mathcal{D}_{1},\dots,\mathcal{D}_{m}\}$ obtained from a reasoning-LLM via the methods described above (i.e., LLM Labeling, Reference Selection, and Generation Reward). The teacher-generated data are used to construct $m*(m-1)/2$ document pairs $(\mathcal{D}_{i},\mathcal{D}_{j})$. Subsequently, an online ranker, which operates as a pointwise ranking LLM, assumes the role of the student model. This student is optimized using the RankNet loss, a pairwise loss function that measures the accuracy of the relative ordering between items, as defined by
@@ 4.3.2 Lightweight Ranking
PLM-based Ranking. In PLM-based ranking systems, traditional end-to-end rankers harness robust PLMs, such as those based on BERT or ERNIE (Li et al., 2025a, 2023c, 2024b, 2023b, 2023e, 2023d), to encode the primary features of each document, including the query, title, and content. Additional signals such as authority, quality, recency, and relevance features are fused through a neural network fusion layer to generate a final rank score (Li et al., 2025b). This end-to-end framework integrates both semantic and heuristic features, but it is limited by the representational power of the base model.
@@ 5.2 Robust RAG System
In practical RAG scenarios, the retriever is not always perfect, and the prompts provided to the
@@
generator often contain noisy documents. However, LLMs that have not undergone further optimization tend to be sensitive to such noise in the prompts, which can degrade the overall performance of the RAG system. To address this issue, the AI search system proposes an adversarial tuning in a multi-agent method ATM (Zhu et al., 2024), which leverages a multi-agent system combined with adversarial tuning to enhance the robustness of the generator. An overview of the ATM is illustrated in Fig. 11.
5.2.1 Adversarial Training
@@
5.2.2 Multi-Agent Training
@@
    The Attacker is trained using Direct Preference Optimization (DPO) (Rafailov et al., 2023) with the loss function defined as
@@
    where $A_{\mathrm{ref}}$ is the un-aligned Attacker (reference model), $A_{\theta}$ is the current Attacker to be optimized. $d_{win}^{\prime}$ and $d_{lose}^{\prime}$ represent a pair of fabrications generated by the Attacker, $win$ denotes the one with a higher $\mathrm{PPL}$-based reward, and vice versa.
@@ 5.3 RAG Task Alignment
To better harness the impressive question-answering capabilities of LLMs within RAG scenarios, it is essential to further align LLMs with the specific requirements of RAG tasks. To address this challenge, this work introduces a preference alignment technique for the RAG scenario (PA-RAG) (Wu et al., 2024), which achieves comprehensive alignment between LLMs and RAG requirements through multi-perspective preference optimization.
@@ 5.3.3 Methodology of PA-RAG
The preference optimization stage adheres to a strict optimization sequence, training the model sequentially and independently on response informativeness, response robustness, and citation quality. This sequence is arranged according to the increasing difficulty of preference optimization, similar to curriculum learning (Bengio et al., 2009). Modifying this order can result in performance degradation.
@@ 5.4 Optimization with User Feedback
Although LLMs demonstrate powerful capabilities in understanding human instructions and generating high-quality responses by pre-training on extensive corpora and subsequently fine-tuning on meticulously designed instruction-following datasets, it is often observed that fine-tuned LLMs produce unexpected and even harmful responses that significantly deviate from human expectations. To address this, most LLMs incorporate alignment strategies that aim to guide the models to meet predetermined dimensions, such as generating harmless, useful, and honest answers. Typically, these alignment methods (Zheng et al., 2023; Touvron et al., 2023) first assign preference signals to answers generated by various models and then employ techniques, such as RLHF or DPO, to train the models based on these preference signals. However, these approaches usually incur high labor and time costs, as well as significant sampling inference overhead. Moreover, some studies (Chang et al., 2023; Bai et al., 2022) have explored the use of AI-assisted annotation or AI-guided feedback as alignment signals to mitigate the reliance on human effort. Nevertheless, the feedback distributions from these sources generally do not align with online user behaviors and tend to be time-varying and heterogeneous. To tackle the above problems, the AI search system first explores the direct alignment of LLMs with online human behavior and proposes an LLM alignment method Reinforcement Learning with Human Behaviors (RLHB). Specifically, RLHB employs a multi-model simultaneous training mechanism with the target LLM as the generator and another auxiliary LLM as the discriminator. They are trained adversarially, ensuring that <query, response, feedback > originates from real online interactions. In inference, the aligned generator accepts the user query along with the most preferred behavior as input signals to generate responses. Compared to RLHF, RLHB eliminates the need for extensive annotation and is adaptable to evolving user behaviors by updating behavior modeling in natural-language form.
[…]
@@
where $\mathcal{M}_{e}$ and $\mathcal{M}_{g}$ denote the expert and generated demonstration sets, respectively. The policy $\pi(a_{t}|s_{t},b_{t})$ is optimized to generate feedback-conditioned responses using the clipped Proximal Policy Optimization (PPO) (Schulman et al., 2017) surrogate objective:
@@
where $\gamma$ is the discount factor, $\lambda$ is the GAE parameter, $V_{\phi}$ is the value function with parameters $\phi$ The reward signal $r_{t}$ includes the discriminator output, regularized with a KL divergence penalty as
@@ 5.5 Multi-Agent Joint Optimization
LLMs have revolutionized various applications, including question answering (Asai et al., 2023), information retrieval (Xiong et al., 2024), and various forms of reasoning (Hao et al., 2023); however, their reliance on static pre-training inhibits timely updates to internal knowledge, often resulting in outdated or fabricated outputs (Zhao et al., 2023b). RAG addresses this limitation by integrating current external knowledge into the generation process, thereby mitigating hallucinations and enhancing response accuracy. Contemporary RAG systems are structured as complex pipelines composed of interdependent modules, such as query rewriting (Jiang et al., 2024), first-stage retrieval (Lewis et al., 2020a), re-ranking (Salemi and Zamani, 2024b), document processing (Ke et al., 2024), and answer generation (Shao et al., 2023b), that are traditionally optimized independently via supervised fine-tuning (SFT) on human-annotated data. This modular approach can lead to a misalignment between individual module objectives (e.g., training retrieval modules to maximize metrics like NDCG (Järvelin and Kekäläinen, 2002)) and the overarching goal of generating accurate answers, since improved relevance does not necessarily yield higher answer quality (Cuconasu et al., 2024). Prior efforts (Guu et al., 2020; Lee et al., 2019; Cuconasu et al., 2024; Izacard and Grave, 2020; Zamani and Bendersky, 2024; Salemi and Zamani, 2024a; Ke et al., 2024) to overcome these challenges have employed end-to-end optimization techniques, including the propagation of rewards from the final answer through attention distributions or generation probabilities, as well as reinforcement learning methods like DPO and PPO; however, these strategies are typically confined to simplified two-module pipelines and fail to capture the intricate interdependencies and cooperative dynamics present in more sophisticated multi-module systems
@@
To address the above limitations, the AI search system proposes a multi-module joint optimization algorithm for RAG (MMOA-RAG(Chen et al., 2025b). By modeling each intermediate module, such as the Planner, Web Search Tool, the Executor, and the Writer, as an independent agent, the optimization process is formulated as a multi-agent collaborative RL task. MMOA-RAG leverages the multi-agent PPO (MAPPO) algorithm (Yu et al., 2022) to align the individual objectives of all modules with the shared goal of maximizing the quality and accuracy of generated answers, as evaluated through metrics like the F1 score against ground-truth responses. MMOA-RAG offers notable flexibility in its choice of reward functions and pipeline configurations, allowing it to accommodate diverse system designs while maintaining the cooperative alignment required for efficient optimization. Unlike earlier approaches, which often focus on optimizing simplified one-module or two-module RAG systems or fail to account for the interdependencies among components, MMOA-RAG provides a robust and generalizable framework for optimizing complex, multi-module RAG pipelines. In contrast to methods based on DPO or single-agent PPO, MMOA-RAG excels at fostering effective collaboration among modules, ensuring that the entire system operates cohesively to produce high-quality responses to knowledge-intensive tasks. Next, we will detail the proposed framework through the following algorithmic description.
@@
As shown in Figure 13, MMOA-RAG consists of four components. Specifically, the Planner refines the initial query $q$, breaking down complex or ambiguous questions into manageable sub-questions $subq$ to improve retrieval efficiency. Web Search Tool then gathers relevant documents for each sub-question, producing a candidate document set $D$. Given the candidate document set, the Executor contained in the Executor then identifies a refined subset $D_{\text{selected}}$ that most effectively supports generating the final answer. Using this filtered subset, the Writer produces the predicted response $Ans_{\text{predict}}$ to the original query. The Planner, Executor, and Writer are implemented using LLMs, enabling them to operate as RL agents that update their parameters based on reward signals. For efficiency, these modules share a single LLM architecture. As Web Search Tool is not modeled as an RL agent due to its unique requirements, it is treated as part of the external environment in reinforcement learning.
@@
Given the inherently cooperative nature of the modules in the RAG pipeline, this shared reward, $R_{\text{shared}}$, serves as a unified signal to guide the optimization of all RL agents. This approach builds on established practices in multi-agent reinforcement learning (MARL) literature (Yu et al., 2022; Rashid et al., 2020; Chen et al., 2022), where shared rewards are leveraged to promote collaboration among agents. To ensure training stability and enhance convergence during the joint optimization process, the framework incorporates penalty mechanisms for individual agents. These penalties refine agent behaviors and discourage inefficient actions described as follows:
@@
[…]
@@
Specifically, for the Planner, public datasets from Rewrite-Retrieve-Read pipelines (Ma et al., 2023) are utilized for SFT by training the module to reformulate an input query $q$ into a series of sub-questions $\text{subq}=\{q_{\text{subq}_{1}},q_{\text{subq}_{2}},\dots\}$, making retrieval more effective.
@@
where $\alpha$ is a hyperparameter balancing the two terms. The Actor loss $\mathcal{L}_{\text{Actor}}$ incorporates advantage estimates $\hat{A}_{t}^{i}$, which is computed using Generalized Advantage Estimation (GAE) (Schulman et al., 2015), and the importance sampling ratio $r_{t}^{i}$ to perform stable policy updates:

### v2

@@ 4.2.1 LLM Labeling
Compared to manual labeling processes, LLMs can efficiently annotate vast datasets at scale, thereby substantially reducing both the temporal and human resource investments required for training retrieval and ranking models. Furthermore, LLMs deliver highly consistent annotations, mitigating the subjective biases and inter-annotator variability that commonly afflict human-generated labels. By adjusting prompt designs or fine-tuning for specific domains or tasks, LLMs exhibit exceptional flexibility in meeting diverse annotation requirements. As previously noted, the AI search system utilizes listwise methods for document annotation, as these approaches effectively strike a balance between effectiveness and efficiency. In this work, the AI search system primarily utilizes two methods as follows.: RankGPT [Sun2023IsCG] and TourRank [Chen2024TourRankUL].
RankGPT. When LLMs come to ranking tasks, especially in the context of passage re-ranking, the limited context window of LLMs poses a significant challenge. Most LLMs have a maximum context length that restricts the number of passages they can process simultaneously. This limitation is particularly problematic in scenarios where a large number of candidate documents need to be ranked, such as in web search or large-scale IR systems. To address this challenge, the AI search system utilizes a strategy called the sliding window approach RankGPT [Sun2023IsCG]. This method allows the efficient re-ranking of large candidate sets without exceeding the context window limitations of the LLM. By breaking down the ranking task into smaller, manageable segments, RankGPT ensures that the LLM can effectively process and rank each subset of documents while maintaining the overall ranking quality.
@@
  • Initialization. Given the retrieved set of $M$ candidate documents from the retrieval model, the sliding window strategy begins with the last $w$ documents, i.e., from the (Mw+1)(M-w+1)-th to the $M$-th passage. These documents are input to the LLM, which returns a ranked order for this subset. The choice of $w$ (window size) is crucial as it determines the number of documents processed in each iteration. A larger window size allows the LLM to consider more documents at once, potentially improving the ranking quality, but also increases the computational load and the risk of exceeding the context window limitations.
  • Sliding. Once the initial window is ranked, the window slides backward by $s$ documents. The new window now covers documents from (Mws+1)(M-w-s+1) to (Ms)(M-s). The LLM is again used to rank this subset, which may overlap with the previous window. The step size $s$ controls the overlap between consecutive windows. A smaller step size increases the overlap, allowing the LLM to compare more documents across windows, but also increases the number of iterations required to cover all documents.
@@
TourRank. LLMs have shown great potential in zero-shot document ranking. However, existing LLM-based ranking methods face several limitations. First, the maximum context length of LLMs restricts the number of documents that can be processed simultaneously, making it challenging to rank a large number of documents at once. Second, the ranking results are highly sensitive to the order in which documents are presented to the LLM, leading to inconsistent outcomes. Third, achieving a balance between computational cost and ranking performance remains a significant challenge. To address these issues, the AI search system introduces a novel approach inspired by the structure of sports tournaments, named TourRank [Chen2024TourRankUL].
@@ 4.2.4 Distillation of LLM Ranking
Consider a query $q$ and its corresponding final ranking of candidate documents $\{\mathcal{D}_{1},\dots,\mathcal{D}_{m}\}$ obtained from a reasoning-LLM via the methods described above (i.e., LLM Labeling, Reference Selection, and Generation Reward). The teacher-generated data are used to construct m(m1)/2m*(m-1)/2 document pairs $(\mathcal{D}_{i},\mathcal{D}_{j})$. Subsequently, an online ranker, which operates as a pointwise ranking LLM, assumes the role of the student model. This student is optimized using the RankNet loss, a pairwise loss function that measures the accuracy of the relative ordering between items, as defined by
@@ 4.3.2 Lightweight Ranking
PLM-based Ranking. In PLM-based ranking systems, traditional end-to-end rankers harness robust PLMs, such as those based on BERT or ERNIE [li2025s, li2025rankelectra, li2023s2phere, DBLP:journals/ml/LiXKBWCY24, DBLP:conf/icdm/0006XKSCWY23, DBLP:journals/tkde/LiXWKLLBWCDY23, DBLP:conf/pkdd/LiXKWSCCY23, li2025rankexpert], to encode the primary features of each document, including the query, title, and content. Additional signals such as authority, quality, recency, and relevance features are fused through a neural network fusion layer to generate a final rank score [li2025m2oerank]. This end-to-end framework integrates both semantic and heuristic features, but it is limited by the representational power of the base model.
@@ 5.2 Robust RAG System
In practical RAG scenarios, the retriever is not always reliable and may retrieve noisy or
@@
irrelevant documents. Consequently, these noisy references may misguide the generator, resulting in inaccurate or less robust outputs from the RAG system. To address this issue, the AI search system proposes an adversarial tuning in a multi-agent method ATM [ATM], which leverages a multi-agent system combined with adversarial tuning to enhance the robustness of the generator. An overview of the ATM is illustrated in Fig. 11.
5.2.1 Adversarial Setups for Robust RAG
@@
5.2.2 Multi-Agent Adversarial Training
@@
    The Attacker is trained using Direct Preference Optimization (DPO) [rafailov2023direct] with the loss function defined as
@@
    where $A_{\mathrm{ref}}$ is the un-aligned Attacker (reference model), $A_{\theta}$ is the current Attacker to be optimized. $d_{win}^{\prime}$ and $d_{lose}^{\prime}$ represent a pair of fabrications generated by the Attacker, winwin denotes the one with a higher $\mathrm{PPL}$-based reward, and vice versa.
@@ 5.3 RAG Task Alignment
To better harness the impressive question-answering capabilities of LLMs within RAG scenarios, it is essential to further align LLMs with the specific requirements of RAG tasks. To address this challenge, this work introduces a preference alignment technique for the RAG scenario (PA-RAG) [PA-RAG], which achieves comprehensive alignment between LLMs and RAG requirements through multi-perspective preference optimization.
@@ 5.3.3 Methodology of PA-RAG
The preference optimization stage adheres to a strict optimization sequence, training the model sequentially and independently on response informativeness, response robustness, and citation quality. This sequence is arranged according to the increasing difficulty of preference optimization, similar to curriculum learning [bengio2009curriculum]. Modifying this order can result in performance degradation.
@@ 5.4 Optimization with User Feedback
Although LLMs demonstrate powerful capabilities in understanding human instructions and generating high-quality responses by pre-training on extensive corpora and subsequently fine-tuning on meticulously designed instruction-following datasets, it is often observed that fine-tuned LLMs produce unexpected and even harmful responses that significantly deviate from human expectations. To address this, most LLMs incorporate alignment strategies that aim to guide the models to meet predetermined dimensions, such as generating harmless, useful, and honest answers. Typically, these alignment methods [zheng2023secrets, touvron2023llama] first assign preference signals to answers generated by various models and then employ techniques, such as RLHF or DPO, to train the models based on these preference signals. However, these approaches usually incur high labor and time costs, as well as significant sampling inference overhead. Moreover, some studies [DBLP:journals/corr/abs-2306-11816, bai2022constitutional] have explored the use of AI-assisted annotation or AI-guided feedback as alignment signals to mitigate the reliance on human effort. Nevertheless, the feedback distributions from these sources generally do not align with online user behaviors and tend to be time-varying and heterogeneous. To tackle the above problems, the AI search system first explores the direct alignment of LLMs with online human behavior and proposes an LLM alignment method Reinforcement Learning with Human Behaviors (RLHB). Specifically, RLHB employs a multi-model simultaneous training mechanism with the target LLM as the generator and another auxiliary LLM as the discriminator. They are trained adversarially, ensuring that <query, response, feedback > originates from real online interactions. In inference, the aligned generator accepts the user query along with the most preferred behavior as input signals to generate responses. Compared to RLHF, RLHB eliminates the need for extensive annotation and is adaptable to evolving user behaviors by updating behavior modeling in natural-language form.
The user behavior used in RLHB includes explicit and implicit behavior. More specifically, explicit behavior refers to users’ proactive feedback behavior, such as the click behavior on Like and Dislike buttons. The implicit behavior usually involves indirect and non-perceived interactions, such as Page View (PV) and Clicks (i.e., the number of times an answer is clicked). In particular, Page View, Clicks, and Like are positive indicators, and Dislike is negative. At the implementation level, these indicators are first smoothed using a logarithmic transformation and then discretized into equal intervals. Subsequently, reward shaping is performed during reinforcement learning by computing $\left(\emph{Like}-\emph{Dislike}\right)/\left(\emph{PV}+\emph{Clicks}\right)$, integrating multiple reward signals into a holistic scalar. Next, we present the details of RLHB as follows.
@@
where $\mathcal{M}_{e}$ and $\mathcal{M}_{g}$ denote the expert and generated demonstration sets, respectively. The policy $\pi(a_{t}|s_{t},b_{t})$ is optimized to generate feedback-conditioned responses using the clipped Proximal Policy Optimization (PPO) [schulman2017proximal] surrogate objective:
@@
where $\gamma$ is the discount factor, $\lambda$ is the GAE parameter, $V_{\phi}$ is the value function with parameters $\phi$ . The reward signal $r_{t}$ includes the discriminator output, regularized with a KL divergence penalty as
@@ 5.5 Multi-Agent Joint Optimization
LLMs have revolutionized various applications, including question answering [asai2023self], information retrieval [xiong2024search], and various forms of reasoning [hao2023reasoning]; however, their reliance on static pre-training inhibits timely updates to internal knowledge, often resulting in outdated or fabricated outputs [zhao2023survey]. RAG addresses this limitation by integrating current external knowledge into the generation process, thereby mitigating hallucinations and enhancing response accuracy. Contemporary RAG systems are structured as complex pipelines composed of interdependent modules, such as query rewriting [jiang2024rag], first-stage retrieval [lewis2020retrieval], re-ranking [salemi2024towards], document processing [ke2024bridging], and answer generation [shao2023enhancing], that are traditionally optimized independently via supervised fine-tuning (SFT) on human-annotated data. This modular approach can lead to a misalignment between individual module objectives (e.g., training retrieval modules to maximize metrics like NDCG [dcg]) and the overarching goal of generating accurate answers, since improved relevance does not necessarily yield higher answer quality [cuconasu2024power]. Prior efforts [guu2020retrieval, lee2019latent, cuconasu2024power, izacard2020distilling, zamani2024stochastic, salemi2024learning, ke2024bridging] to overcome these challenges have employed end-to-end optimization techniques, including the propagation of rewards from the final answer through attention distributions or generation probabilities, as well as reinforcement learning methods like DPO and PPO; however, these strategies are typically confined to simplified two-module pipelines and fail to capture the intricate interdependencies and cooperative dynamics present in more sophisticated multi-module systems
@@
To address the above limitations, the AI search system proposes a multi-module joint optimization algorithm for RAG (MMOA-RAG) [chen2025improving]. By modeling each intermediate module, such as the Planner, Web Search Tool, the Executor, and the Writer, as an independent agent, the optimization process is formulated as a multi-agent collaborative RL task. MMOA-RAG leverages the multi-agent PPO (MAPPO) algorithm [yu2022surprising] to align the individual objectives of all modules with the shared goal of maximizing the quality and accuracy of generated answers, as evaluated through metrics like the F1 score against ground-truth responses. MMOA-RAG offers notable flexibility in its choice of reward functions and pipeline configurations, allowing it to accommodate diverse system designs while maintaining the cooperative alignment required for efficient optimization. Unlike earlier approaches, which often focus on optimizing simplified one-module or two-module RAG systems or fail to account for the interdependencies among components, MMOA-RAG provides a robust and generalizable framework for optimizing complex, multi-module RAG pipelines. In contrast to methods based on DPO or single-agent PPO, MMOA-RAG excels at fostering effective collaboration among modules, ensuring that the entire system operates cohesively to produce high-quality responses to knowledge-intensive tasks. Next, we will detail the proposed framework through the following algorithmic description.
@@
As shown in Figure 13, MMOA-RAG consists of four components. Specifically, the Planner refines the initial query $q$, breaking down complex or ambiguous questions into manageable sub-questions subqsubq to improve retrieval efficiency. Web Search Tool then gathers relevant documents for each sub-question, producing a candidate document set $D$. Given the candidate document set, the Executor contained in the Executor then identifies a refined subset $D_{\text{selected}}$ that most effectively supports generating the final answer. Using this filtered subset, the Writer produces the predicted response $Ans_{\text{predict}}$ to the original query. The Planner, Executor, and Writer are implemented using LLMs, enabling them to operate as RL agents that update their parameters based on reward signals. For efficiency, these modules share a single LLM architecture. As Web Search Tool is not modeled as an RL agent due to its unique requirements, it is treated as part of the external environment in reinforcement learning.
@@
Given the inherently cooperative nature of the modules in the RAG pipeline, this shared reward, $R_{\text{shared}}$, serves as a unified signal to guide the optimization of all RL agents. This approach builds on established practices in multi-agent reinforcement learning (MARL) literature [yu2022surprising, rashid2020monotonic, chen2022ptde], where shared rewards are leveraged to promote collaboration among agents. To ensure training stability and enhance convergence during the joint optimization process, the framework incorporates penalty mechanisms for individual agents. These penalties refine agent behaviors and discourage inefficient actions described as follows:
@@
Moreover, the observation of the Executor can be represented as $O_{\text{S}}=\{Prompts,q,D\}$, where $Prompts$ are guiding instructions, $q$ is the query, and $D$ is a set of $K$ candidate documents. Its action space $A_{\text{S}}=\{\text{"0"},\text{"1"},...,\text{"K-1"},\text{"Document"},\text{","}\}$ is constrained to document selection, reducing the exploration space for efficient training. The reward function $R_{\text{S}}$ incorporates the penalty $P_{\text{S}}$, which is set to 1-1 when the selected document IDs are invalid (e.g., duplicates or incorrect formats) and 0 otherwise. $R_{\text{S}}$ can be represented as follows:
@@
Specifically, for the Planner, public datasets from Rewrite-Retrieve-Read pipelines [ma2023query] are utilized for SFT by training the module to reformulate an input query $q$ into a series of sub-questions $\text{subq}=\{q_{\text{subq}_{1}},q_{\text{subq}_{2}},\dots\}$, making retrieval more effective.
@@
where $\alpha$ is a hyperparameter balancing the two terms. The Actor loss $\mathcal{L}_{\text{Actor}}$ incorporates advantage estimates $\hat{A}_{t}^{i}$, which is computed using Generalized Advantage Estimation (GAE) [schulman2015high], and the importance sampling ratio $r_{t}^{i}$ to perform stable policy updates: